McClain, Inc. purchased a group of assets for $3,000. The assets purchased (and their market values) were land ($1,479), building ($1,530), equipment ($816), and a truck ($1,275). McClain allocates the cost of these assets based on the relative fair market values. How much of the cost is allocated to the equipment? Round calculation to 2 decimals … please pay attention … 0.3597823 rounds to 36%; Your final answer should be in whole dollars (NO decimals)
Answer #1

Total fair value = 1479+1530+816+1275 = 5100

Cost allocated to equipment

= (816/5100)*3000

= 0.16*3000

= 480